# SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

### v1

@@ 3.2 Layer-wise Bit Allocation
We formulate layer-wise bit allocation as a discrete optimization problem. Let $n$ denote the total number of layers, and let $B=\{b_{1},\dots,b_{K}\}$ be the set of allowed bit-widths. For each layer $i\in\{1,\dots,n\}$ and bit-width $b\in B$, we introduce a binary variable $I_{i,b}\in\{0,1\}$, where $I_{i,b}=1$ indicates that layer $i$ is assigned bits $b$.
@@
Let $\Delta L_{i}(b)$ denote the delta loss when quantizing layer $i$ to $b$ bits. Given a target average bit-width $T$, the optimization problem can be written as:
@@
This optimization problem can be solved using dynamic programming  (bellman1966dynamic) or as an Integer Linear Program  (wolsey2020integer; bertsimas1997introduction), both of which are well-established methods. We omit the details here and refer the reader to the existing literature.
3.3 Quantization Parameters Initialization
In SignRoundV1 (cheng2024optimize), all quantization parameters are initialized with trivial values, i.e., weight clipping is set to 1.0 and the rounding perturbation is 0. A better initialization could potentially improve performance, similar to the effect of careful initialization in training neural networks  (he2015delving; glorot2010understanding). Although directly initializing the rounding values is challenging, prior works (lin2024awq; badri2023hqq; ggerganov_llamacpp) have proposed training-free methods to search for weight clipping by aligning layer outputs or optimizing certain objectives. While effective, these approaches require layer-wise forward passes, which can be slow, or rely on floating-point zero points for better accuracy.
Inspired primarily by the importance matrix in llama.cpp (ggerganov_llamacpp), we adopt a simple variant to initialize the scale. The search process is formulated as
$\min_{s\in S}\;\frac{1}{N}\sum_{i=1}^{N}\left((W_{F}-W_{q})\circ max(\bar{A})^{2}\right)_{i}^{2}$ (9)
where $S$ denotes a predefined set of candidate scales, and $max(\bar{A})$ represents the input channel-wise maximum absolute values, calibrated from a set of samples.
For symmetric quantization, which is the default setting used in this work, the set $S(s_{0},...,s_{c})$ is constructed from the following candidates
$s_{i}=\frac{\max(\mathbf{|W|})}{2^{\text{bit}-1}+\epsilon_{i}}.$ (10)
with $\epsilon_{i}$ sampled from
$\epsilon\in\mathrm{range}(-0.9,\,0.9,\,\text{step}=0.01)$ (11)
Following SignRoundV1 (cheng2024optimize), the best scale $s_{init}$ that minimizes Eq. 9 is further refined by a learnable parameter $\alpha$, which we constrain to the range [0.5, 1.5] in our experiments.
$s=s_{init}*\alpha$ (12)
3.4 Hyperparameters and Tricks
For DeltaLoss, we use only 16 calibration samples with a sequence length of 256. For tuning, we follow the setup in SignRoundV1 (cheng2024optimize). Each transformer block is optimized for 200 steps using sign gradient descent, with a learning rate of $1/steps$ and a batch size of 8. The sequence length is fixed at 2048. To reduce quantization cost, we lower the default number of calibration samples from 512 to 128. We additionally provide a higher-cost recipe for improved accuracy, denoted as Ours*, where the number of steps is increased from 200 to 500, the learning rate is set to $2.0/steps$ and the number of calibration samples is restored to 512. Automatic mixed precision (AMP) is applied throughout to improve tuning efficiency.
@@ 3.4 Hyperparameters and Tricks
$L_{\text{eff}}=\sum_{L_{i}\in\mathcal{L}\setminus\text{TopK}(\mathcal{L},k)}L_{i}.$ (13)
Table 1: Average accuracies across five tasks (Section 4.2) using W2A16 at 2 bits and mixed W2A16 and W4A16G128 at 2.5 bits. "G" denotes the group size, and "Avg. Bits" denotes the average per-weight bit-width. Ours* indicates our improved recipe (Section 3.4). ’-’ denotes the missing data from the original work.
@@
Ours* 64 60.28 64.57 70.60 63.56 72.32
Table 2: The average accuracies and recovery rates (%) across 10 tasks(detailed in Section 4.3) with MXFP4 at 4bits and mixed MXFP4/8 for others. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning. The "I" in the model name is short for "Instruct".

### v2

@@ 3.2 Layer-wise Bit Allocation
We formulate layer-wise bit allocation as a discrete optimization problem. Let $n$ denote the total number of layers, and let $B=\{b_{1},\dots,b_{K}\}$ be the set of allowed bit-widths. For each layer $i\in\{1,\dots,n\}$ and bit-width $b\in B$, we introduce a binary variable $I_{i,b}\in\{0,1\}$, where $I_{i,b}=1$ indicates that layer $i$ is assigned bit-width $b$.
@@
Let $\Delta L_{i}(b)$ denote the DeltaLoss when quantizing layer $i$ to $b$ bits. Given a target average bit-width $T$, the mixed-precision assignment can be written as:
@@
where $P_{i}$ denotes the number of parameters in layer ii. This optimization problem can be solved via dynamic programming [5] or as an integer linear program [67, 6], both of which are well-established approaches. We omit the details here and refer the reader to the existing literature.
3.3 Tuning Quantization Parameters
In SignRoundV1 [12], all quantization parameters are initialized with trivial values, specifically, the weight clipping parameters $\alpha$ and $\beta$ are set to 1.0, and the rounding perturbation $v$ is initialized to 0. A better initialization could potentially improve performance, similar to the effect of careful initialization in training neural networks [28, 26]. Although initializing the rounding values is challenging, prior works [39, 3, 25] have proposed training-free methods to search for weight clipping by aligning layer outputs or optimizing certain objectives. While effective, these approaches require layer-wise forward passes, which can be slow, or rely on floating-point zero points for better accuracy.
Inspired primarily by the importance matrix in llama.cpp [25], we adopt a simple variant to initialize the scale. The search process is formulated as
$\min_{s\in S}\;\frac{1}{N}\sum_{i=1}^{N}\left((W_{\mathrm{f}}-W_{\mathrm{q}})\odot\overline{A}\right)_{i}^{2},$ (10)
where $S$ denotes a predefined set of candidate scales, and $\overline{A}$ represents the channel-wise average of input activations, calibrated from a set of samples.
For symmetric quantization, which is the default setting used in this work, the set $S=\{s_{0},\ldots,s_{c}\}$ is constructed from the following candidates:
$s_{i}=\frac{\max(|\mathbf{W}|)}{2^{b-1}+\epsilon_{i}},\quad\epsilon_{i}\in\{-t,-t+\delta,\ldots,t\}.$ (11)
$t$ and $\delta$ are fixed for each scheme. Specifically, we set $t=0.9$ and $\delta=0.01$ for W2A16. We refine the optimal scale $s_{init}$ with a learnable parameter $\alpha$, i.e., $s=s_{init}\cdot\alpha$, where $\alpha\in[0,2.0]$. Finally, we follow SignRoundV1 [12] to tune $v$ in Equation 3.
3.4 Tuning Stabilization via Loss Filtering
@@
$L_{\text{eff}}=\frac{1}{|\mathcal{L}\setminus\mathrm{TopK}(\mathcal{L},k)|}\sum_{L_{i}\in\mathcal{L}\setminus\mathrm{TopK}(\mathcal{L},k)}L_{i}.$ (12)
3.5 Hyperparameters
For DeltaLoss, we use only 16 calibration samples with a sequence length of 256. For tuning, we follow the setup in SignRoundV1 [12]. Each transformer block is optimized for 200 steps using signed gradient descent, with a learning rate of $1/\text{steps}$ and a batch size of 8. The sequence length is fixed at 2048. To reduce quantization cost, we lower the default number of calibration samples from 512 to 128. We additionally provide a higher-cost recipe for improved accuracy, denoted as Ours*, where the number of steps is increased from 200 to 500, the learning rate is set to $2.0/\text{steps}$ and the number of calibration samples is restored to 512. Automatic mixed precision (AMP) is applied throughout to improve tuning efficiency.
Table 1: Average accuracies across five tasks (Section 4.2) using W2A16 at 2 bits and mixed W2A16 and W4A16G128 at 2.5 bits. “G” denotes the group size, and “Avg. Bits” denotes the average per-weight bit-width. Ours* indicates our improved recipe (Section 3.5). “-” denotes missing data from the original work.
@@ 3.5 Hyperparameters
Ours* 64 60.44 64.57 70.60 63.56 72.32
4 Experiments
4.1 Experimental Setup
Models and Benchmarks.
We evaluate SignRoundV2 on two prominent LLM families: LLaMA [61, 27] and Qwen [4, 70]. Our evaluation spans both low-bit and mixed-precision quantization settings across a comprehensive suite of standard benchmarks. All evaluations are conducted in a zero-shot setting using the lm-evaluation-harness framework [24]. To ensure fair comparisons, all methods within each experimental group are evaluated with a version of lm_eval that produces identical or closely aligned baseline results, unless explicitly stated otherwise. The benchmark suite includes ARC-Challenge [15], ARC-Easy [15], BoolQ [14], HellaSwag [74], LAMBADA [55], MMLU [29], OpenBookQA [50], PIQA [7], TruthfulQA [40], and WinoGrande [56].
Quantization Configurations.
Following SignRoundV1 [12], we utilize the Pile [23] as the calibration dataset. Unlike SignRoundV1, we adopt symmetric quantization to enhance hardware and kernel compatibility. All experiments are executed on NVIDIA A100 (80GB) GPUs. For clarity, the bit costs associated with scales and zero points are omitted unless otherwise specified. For MXFP4 and MXFP8 settings, we strictly adhere to the standard definitions specified in [1].
4.2 Comparison of Weight-Only Quantization Methods
Table 1 reports average accuracies across five representative tasks for W2A16 with and without mixed-precision support. In uniform W2A16 settings, SignRoundV2 consistently outperforms established PTQ baselines, including GPTQ [22], AWQ [39], OmniQuant [57], and SignRoundV1 [12], by a substantial margin. Compared to high-complexity methods such as AQLM [20], QuIP# [63], and EfficientQAT [10], our approach achieves competitive performance on large-scale models while maintaining a significantly lower optimization overhead. Furthermore, at slightly higher average bit-widths (e.g., 2.5 bits), our adaptive layer-wise strategy effectively preserves model quality, demonstrating robust recovery capabilities in extreme compression regimes. Detailed INT2/4 mixed-bit results are in Appendix E.
4.3 Comparison of Adaptive Bit-Width Quantization
Table 2 shows that SignRoundV2 achieves strong accuracy across evaluated bit-widths, with particularly large gains in the ultra-low-bit regime and competitive performance at higher bit-widths. Following SFMP [53], we report average accuracy across six tasks, prioritizing acc_norm. To ensure fair comparison, we account for scaling factor and zero-point overhead while restricting the bit-width search space to a maximum difference of 1. For reference, we provide the accuracy of uniform-bit SignRoundV2 at 3 and 4 bits (shown in gray), serving as non-mixed baselines. Notably, SFMP [53] and AMQ [35] use AWQ [39] as their base weight quantizer. Against these baselines, SignRoundV2 achieves significant gains in ultra-low-bit regimes; e.g., on Llama-3.1-8B at 2.5 bits, it improves the recovery rate over AMQ by 12.24 percentage points. SignRoundV2 achieves competitive performance with SFMP while maintaining hardware-friendly per-tensor granularity, avoiding kernel-level overhead inherent in SFMP’s per-group strategy. These results, supported by Table 6, highlight our favorable accuracy–cost trade-off. We further compare DeltaLoss against naive strategies such as head/tail fallback in Appendix A.
4.4 Performance under MXFP Quantization
Table 3 presents our results using the MX format [1]. Given that most existing PTQ methods do not support MX formats or lack comparable baselines, we primarily compare SignRoundV2 against SignRoundV1 [12] and RTN. SignRoundV2 exhibits performance gains in the 4-bit regime, surpassing RTN and outperforming SignRoundV1 in most scenarios. At average bit-widths of 4.5 and 5 bits, our method approaches full-precision performance on most evaluated models. Detailed results for MXFP4/8 are deferred to Appendix C.2, while Appendix D reports the NVFP4 results.
Table 2: Comparison of mixed-bit algorithms. Average accuracy (acc_norm preferred) and recovery rate across six tasks (see Appendix B for details). Gray rows: uniform 3-bit and 4-bit baselines quantized with SignRoundV2. Note: Ours employs hardware-friendly per-tensor mixed-bit granularity, avoiding per-group granularity overhead in SFMP.
Avg. bits Method Llama3.1-8B Qwen3-8B Qwen3-14B Qwen3-32B
16 Ours 75.83 74.07 77.41 78.05
SFMP 75.01 74.20 77.38 77.99
2.5 AMQ 58.65 (78.19%) 58.62 (79.00%) 66.56 (86.02%) 69.37 (88.95%)
SFMP 64.34 (85.78%) 66.16 (89.16%) 71.12 (91.91%) 73.33 (94.02%)
Ours 68.57 (90.43%) 66.82 (90.20%) 72.63 (93.83%) 74.96 (96.04%)
3 AMQ 68.78 (91.69%) 68.40 (92.18%) 71.32 (92.17%) 74.02 (94.91%)
SFMP 69.74 (92.89%) 71.36 (96.17%) 75.30 (97.31%) 76.24 (97.76%)
Ours 71.64 (94.48%) 72.21 (97.49%) 74.66 (96.45%) 76.61 (98.16%)
W3G128 72.84 (96.06%) 73.24 (98.88%) 76.39 (98.68%) 76.66 (98.22%)
3.5 AMQ 72.56 (96.73%) 71.65 (96.56%) 75.76 (97.91%) 76.01 (97.46%)
SFMP 72.97 (97.28%) 72.74 (98.03%) 76.89 (99.37%) 77.24 (99.04%)
Ours 73.94 (97.51%) 72.99 (98.54%) 76.32 (98.59%) 77.36 (99.12%)
4 AMQ 73.46 (97.93%) 72.64 (97.90%) 76.62 (99.02%) 77.15 (98.92%)
SFMP 74.33 (99.09%) 73.29 (98.77%) 77.13 (99.68%) 77.89 (99.87%)
Ours 75.07 (99.00%) 73.76 (99.58%) 77.05 (99.53%) 77.48 (99.27%)
W4G128 75.06 (98.99%) 74.11 (100.05%) 77.33 (99.90%) 77.75 (99.62%)
Table 3: Mixed-precision MXFP 4/6-bit quantization results. Average accuracies and recovery rates (%) across 10 tasks. “Avg. Bits”: average bit-width per weight; “DL”: DeltaLoss without tuning; “I”: Instruct models. See Appendix C.1 for details.